# On obstructions to the Euler system method for Rankin-Selberg convolutions

### v2

@@ 4 Examples and counter-examples
Thus either $\alpha u,\alpha u^{-1}$ both lie in ${\mathbb{Z}}_{p}^{\times}$, or they are Galois conjugates of each other lying in some quadratic extension of ${\mathbb{Q}}_{p}$. Since $p$ is odd and prime to the order of $u$, and since $\alpha^{2}\in{\mathbb{Z}}_{p}^{\times}$, $\alpha u,\alpha u^{-1}$ are contained in the maximal unramified extension of ${\mathbb{Q}}_{p}$, so the absolute arithmetic Frobenius $\varphi$ of ${\mathbb{Q}}_{p}^{unr}$ either fixes them both or exchanges them. When $\alpha\notin{\mathbb{Q}}_{p}$, since $\alpha^{2}\in{\mathbb{Q}}$, one has $\varphi(\alpha)=-\alpha$, whence $\varphi(\alpha u)=-\alpha u^{p}$. We thus reach a contradiction if $-1\notin\{u^{p-1},u^{p+1}\}$. When $\alpha\in{\mathbb{Q}}_{p}$, one has $\varphi(\alpha u)=\alpha u^{p}$, so we reach a contradiction if $1\notin\{u^{p-1},u^{p+1}\}$.
Assuming that $\varepsilon_{f}$ is the trivial character, that $F={\mathbb{Q}}$, and that $\mathfrak{p}$ is coprime to $2|\rho_{g}(G_{{\mathbb{Q}}})|$, we have thus proved the following. Pick, for every coset $C$ of $G_{H(\mu_{p^{\infty}})}$ in $G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$, some $\alpha_{C}\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ given by Corollary 2, and let $V_{C}\subset L_{0}$ be a set containing the eigenvalues of non-scalar matrices $N\in\rho_{g}(C)$ such that $\det{N}\neq 1$. If, for each $C$ such that $\alpha_{C}\notin{\mathbb{Z}}_{p}^{\times}$, $V_{C}$ does not contain an element $u$ such that $-1\in\{u^{p-1},u^{p+1}\}$, and, for each $C$ such that $\alpha_{C}\in{\mathbb{Z}}_{p}^{\times}$, $V_{C}$ does not contain an element such that $1\in\{u^{p-1},u^{p+1}\}$, then $T_{f,g,\mathfrak{p}}$ does not satisfy (wE).
@@
Let $D$ be the derived subgroup of $\rho_{g}(G_{{\mathbb{Q}}})$ and $Z\leq\rho_{G}(G_{{\mathbb{Q}}})$ be the subgroup of scalar matrices. Since $g$ is of $A_{5}$-type, $D\subset\operatorname{SL}_{2}(L_{0})$ is a finite group whose projective image is $A_{5}$. In particular, it has even order, hence contains $-I_{2}$. Thus we have an exact sequence $1\rightarrow\{\pm I_{2}\}\rightarrow D\rightarrow A_{5}\rightarrow 1$. Moreover, $D$ and $\rho_{g}(G_{{\mathbb{Q}}})$ have the same projective image, so $\rho_{g}(G_{{\mathbb{Q}}})=ZD$. Since $\det{\rho_{g}(G_{{\mathbb{Q}}})}=\{\pm 1\}$ and $D\subset\operatorname{SL}_{2}(L_{0})$, one has $Z=\{\pm iI_{2},\pm I_{2}\}$, so that $\rho_{g}(G_{{\mathbb{Q}}})=D\cup iD$. Since $D\subset\rho_{g}(G_{H})\subset\rho_{g}(G_{{\mathbb{Q}}})$ and $\varepsilon_{g}(G_{H})=1$, one has $\rho_{g}(G_{H})=D$, and $\rho_{g}(G_{{\mathbb{Q}}}\backslash G_{H})=iD$.
Because $\varepsilon_{g}(G_{H})=\{1\}$, it is enough to test the coset $G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\backslash G_{H(\mu_{p^{\infty}})}$. The element $\alpha\in L^{\times}$ attached to this coset by Corollary 2 is $\alpha=\sqrt{2}$. Because $p\equiv 3,5\pmod{8}$, $\alpha\notin{\mathbb{Q}}_{p}$. Hence, we need to prove that if $u$ is an eigenvalue of any non-scalar matrix in $iD$, one does not have $-1\in\{u^{p-1},u^{p+1}\}$. Let $N\in iD$ be a non-scalar matrix. Then $\det{N}=-1$, and its projective image $\tilde{N}$ has order $2$, $3$ or $5$.
    if $\tilde{N}$ has order $2$, then $N^{2}$ is scalar, so Cayley–Hamilton implies that $N^{2}=I_{2}$, so the eigenvalues of $N$ are ±$\pm 1$, so $u^{p\pm 1}=1\neq-1$.
    if $\tilde{N}$ has order $3$, then $N^{3}$ is scalar. Since $\det{N}=-1$, the eigenvalues of $N$ are $\{ij,ij^{2}\}$ or $\{-ij,-ij^{2}\}$. Either way, any eigenvalue $u$ of $N$ has order $12$, and $p\equiv\pm 1\pmod{12}$, so $-1\notin\{u^{p-1},u^{p+1}\}$.
    if $\tilde{N}$ has order $5$, then $N^{5}$ is scalar. Since $\det{N}=-1$, any eigenvalue $u$ of $N$ is a root of unity of order $20$. Since $p\not\equiv\pm 1\pmod{5}$, the order of $u^{p\pm 1}$ is divisible by $5$, so $-1\notin\{u^{p-1},u^{p+1}\}$ and we are done.
@@
Let $D=\rho_{g}(G_{{\mathbb{Q}}})^{\prime}$ and $Z$ denote the group of scalar matrices contained in $\rho_{g}(G_{{\mathbb{Q}}})$. One has $\det{Z}\subset\varepsilon_{g}(G_{{\mathbb{Q}}})=\{\pm 1\}$ so $Z\subset\{\pm I_{2},\pm iI_{2}\}$. Moreover, the coefficient field of $g$ is ${\mathbb{Q}}(\sqrt{-2})$, so that $g$ has no coefficient equal to ±$\pm 2i$, whence $Z\subset\{\pm I_{2}\}$. As in the previous case, $D\subset\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})$ and its projective image is $S_{4}^{\prime}=A_{4}$, so $D$ has even order, hence $-I_{2}\in D$ and $Z=Z\cap D=\{\pm I_{2}\}$. Thus $D$ is an index two subgroup of $\rho_{g}(G_{{\mathbb{Q}}})$; since $\varepsilon_{g}(G_{H})=\{1\}$, one furthermore has $D\subset\rho_{g}(G_{H})\subset\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})$: thus $D=\rho_{g}(G_{H})=\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})$ and $\rho_{g}(G_{{\mathbb{Q}}}\backslash G_{H})=\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{GL}_{2}(L_{0})_{\det=-1}$.
As above, $\varepsilon_{g}(G_{H})=\{1\}$, so it is enough to test the coset $C:=G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\backslash G_{H(\mu_{p^{\infty}})}$. The element $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ attached to $C$ by Corollary 2 is $\sqrt{6}$. By the previous paragraph, the matrices of $\rho_{g}(C)$ have determinant $-1$ and their projective images lie in $S_{4}\backslash A_{4}$, so their orders are $2$ or $4$. Hence, for any non-scalar $M\in\rho_{g}(C)$, the set of eigenvalues of $M$ is either $\{1,-1\}$ or $\{\omega,-\omega^{-1}\}$ for some primitive $8$-th root of unity $\omega$.
Because $p\equiv\pm 1\pmod{8}$ and $p\not\equiv\pm 1\pmod{12}$, $2$ is a square in ${\mathbb{Z}}_{p}^{\times}$ but $3$ is not, so $\alpha\notin{\mathbb{Z}}_{p}^{\times}$. Thus, we need to prove that $-1$ is not contained in $\{1^{p\pm 1},(-1)^{p\pm 1},\omega^{p\pm 1}\}=\{1,\omega^{p\pm 1}\}$ where $\omega$ is a primitive $8$-th root of unity. Since $p\equiv\pm 1\pmod{8}$, $p\pm 1$ is congruent modulo $8$ to $0,2,-2$, so $\{\omega^{p\pm 1}\}\subset\{1,\omega^{2},\omega^{-2}\}=\{1,\pm i\}$ does not contain $-1$, and we are done.
(c) According to the LMFDB, $L,F,H$ and $\varepsilon_{f},\varepsilon_{g}$ are as above (so $\varepsilon_{g}$ is the quadratic character of $H/{\mathbb{Q}}$), and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\simeq A_{4}$, while $\rho_{g}(G_{{\mathbb{Q}}})$ has cardinality $48$: hence the scalar matrices contained in $\rho_{g}(G_{{\mathbb{Q}}})$ are exactly {±$\{\pm iI_{2},\pm I_{2}\}$. Now, $\rho_{g}(G_{H})$ is a subgroup of index at most $2$ of $\rho_{g}(G_{{\mathbb{Q}}})$; since it is contained in $\operatorname{SL}_{2}(L_{0})$, it does not contain $iI_{2}$, and $\rho_{g}(G_{{\mathbb{Q}}}\backslash G_{H})=i\rho_{g}(G_{H})$. In particular, $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$.
Since $\varepsilon_{g}(G_{H})=\{1\}$, it is enough to consider the coset $G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\backslash G_{H(\mu_{p^{\infty}})}$, to which the element $\alpha$ attached by Corollary 2 is $\sqrt{6}$. Since $p\not\equiv\pm 1\pmod{8}$ and $p\equiv\pm 1\pmod{12}$, $2$ is not a square in ${\mathbb{Z}}_{p}^{\times}$ and $3$ is, so $\alpha\notin{\mathbb{Q}}_{p}$.
Hence we need to show that for any non-scalar matrix $M\in\rho_{g}(G_{{\mathbb{Q}}}\backslash G_{H})=i\rho_{g}(G_{H})$, for any eigenvalue $u$ of $M$, one has $-1\notin\{u^{p-1},u^{p+1}\}$.
@@
    If $\tilde{N}$ has order $3$, then as in case (a), any eigenvalue $u$ of $N$ has order $12$. Since $p\equiv\pm 1\pmod{12}$, one among $u^{p-1}$ and $u^{p+1}$ is equal to $1$, and the other has order $6$, so neither equals $-1$ and we are done.
@@
Because $\varepsilon_{g}(G_{H})=\{1\}$, we only need to consider the coset $C:=G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\backslash G_{H(\mu_{p^{\infty}})}$. An $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ attached to $C$ by Corollary 2 is $\alpha=\sqrt{3}$. Since $p\equiv 5,7\pmod{12}$, $3$ is not a square in ${\mathbb{Q}}_{p}$, so we need to show that if $u$ is any eigenvalue of some non-scalar matrix in $\rho_{g}(C)=\rho_{g}(G_{{\mathbb{Q}}})\backslash\operatorname{SL}_{2}(L_{0})=\rho_{g}(G_{{\mathbb{Q}}})\backslash\rho_{g}(G_{H})$, then $-1\notin\{u^{p-1},u^{p+1}\}$.
@@
    If $\det{N}=-1$ and the projective image of $N$ has order $2$, its eigenvalues are ±$\pm 1$; one has $11^{\pm 1}=(-1)^{p\pm 1}=1\neq-$ 1, so we are done.
    If $\det{N}=-1$ and the projective image of $N$ has order $6$, and $u$ is an eigenvalue of $N$, then $N\sim\Delta_{u,-u^{-1}}$, $u^{6}=u^{-6}$, $u^{3}\neq-u^{-3}$, $u^{2}\neq u^{-2}$. Hence $u^{2}\neq 1$, $u^{6}=\pm 1$ and $u^{6}\neq-1$, so $u^{6}=1$. Therefore, $u^{3(p\pm 1)}=1$, so $-1\notin\{u^{p\pm 1}\}$.
(e) In this case, $L={\mathbb{Q}}(\sqrt{7}),F={\mathbb{Q}}$ and $H={\mathbb{Q}}(\sqrt{-3})$; $g$ has complex multiplication by ${\mathbb{Q}}(\sqrt{-51})$, $\varepsilon_{g}$ is a character of conductor $51$ and order $4$, and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\simeq S_{3}$; $\varepsilon_{f}$ is the trivial character. Since $\varepsilon_{g}(2893)=\pm i$, $\varepsilon_{g}(G_{H})=\{\pm 1,\pm i\}=\varepsilon_{g}(G_{{\mathbb{Q}}})$. The LMFDB shows that $\rho_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $S_{3}\times{\mathbb{Z}}/8{\mathbb{Z}}$. Its center (which is scalar, by Schur’s lemma) is thus the set of $8$-th roots of unity, and $\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is onto with kernel {±$\{\pm I_{2}\}$.
The LMFDB data stipulates that $\ker{\tilde{\rho}_{g}}\not\subset G_{H}$, so $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$. Using MAGMA [1], we can also see that $\ker{\rho_{g}}\subset G_{H}$, so that $[\rho_{g}(G_{{\mathbb{Q}}}):\rho_{g}(G_{H})]=2$.
@@
Since $\rho_{g}(G_{H})$ and $\rho_{g}(G_{{\mathbb{Q}}})$ have the same determinant and projective image, we can split elements of $\rho_{g}(G_{{\mathbb{Q}}})$ in six categories: {±$\{\pm iI_{2},\pm I_{2}\}$, scalar matrices of order $8$, matrices in $C$ with determinant ±$\pm 1$, matrices in $C$ with determinant ±$\pm i$, matrices outside $C$ with determinant ±$\pm 1$, matrices outside $C$ with determinant ±$\pm i$. Then $\rho_{g}(G_{H})$ is exactly made with the matrices of the first, third, and last categories.
An element $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ attached to any $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}\backslash G_{H}$ (resp. any $\sigma\in G_{H(\mu_{p^{\infty}})}$) by Corollary 2 is $\alpha=\sqrt{7}$ (resp. $\alpha=1$). Note that because $p\equiv\pm 5,\pm 43,\pm 67\pmod{168}$, $\sqrt{7}\notin{\mathbb{Q}}_{p}$.
@@
    If $u$ is any eigenvalue of any matrix $N$ in the third or last category such that $\det{N}\neq 1$, then $p$ is not congruent to ±$\pm 1$ modulo the order of $u$.
    If $u$ is any eigenvalue of any matrix $N$ in the fourth or fifth category such that $\det{N}\neq 1$, then $-1\notin\{u^{p-1},u^{p+1}\}$.
Let $N\in C$ be in the third category with $\det{N}\neq 1$. Then $\det{N}=-1$ and $N^{3}$ is scalar with determinant $-1$, so as in (a) any eigenvalue $u$ of $N$ has order $12$. We are done since $p\not\equiv\pm 1\pmod{12}$.
@@
Let $N\in C$ be a matrix in the fourth category. Then $\det{N}=\pm i$ and $N^{3}$ is scalar. Then $N\sim\omega\Delta_{j,j^{2}}$ for some primitive $8$-th root of unity $\omega$, so any eigenvalue $u$ of $N$ has order $24$. Since $p\not\equiv\pm 1\pmod{12}$, none of $u^{2(p\pm 1)}\neq 1$ so $-1\notin\{u^{p\pm 1}\}$, and we are done.
Let $N$ be a matrix in the fifth category such that $\det{N}\neq 1$. Then $\det{N}=-1$ and the projective image of $N$ has order $2$, so the eigenvalues of $N$ are ±$\pm 1$, $11^{p\pm 1}=(-1)^{p\pm 1}=$ 1, so we are done.
@@
The proposition below, on the other hand, shows that, even in the case where the assumption of Proposition 2 is satisfied (and therefore the reasoning Corollary 2 does not apply), the representation can still have Euler-adapted Galois image.
@@
    $[\rho(G_{{\mathbb{Q}}}):\rho(G_{H})]=[\tilde{\rho}_{g}(G_{{\mathbb{Q}}}):\tilde{\rho}_{g}(G_{H})]=2$,
@@
Proof. –  If $\rho_{g}(G_{H})$ contains a matrix with eigenvalues $1$ and $-1$, we are done by Corollary 2. We may thus assume that $\rho_{g}(G_{H})$ does not contain any matrix whose eigenvalues are $1$ and $-1$.
Suppose that $\varepsilon_{g}(G_{HK})=\{\pm 1,\pm i\}$. Then $\varepsilon_{g}(G_{H}\backslash G_{K})$ is also equal to {±$\{\pm 1,\pm i\}$, so we can find $\sigma\in G_{H}\backslash G_{K}$ such that $\varepsilon_{g}(\sigma)=-1$. Because $g$ has CM by $K$, $\rho_{g}(\sigma)$ has trace zero, so by Cayley–Hamilton $\rho_{g}(\sigma)$ is similar to $\Delta_{1,-1}$, which contradicts our assumption.
@@
Since $G_{H}/G_{HK}$ is a group of order $2$, it follows that $\varepsilon_{g}(G_{HK})=\{\pm 1\}$, so $\varepsilon_{g}(G_{H}\backslash G_{K})=\{\pm i\}$. Let $H_{0}\subset H$ be the subfield such that $\ker{\varepsilon_{f}}=G_{H_{0}}$. Since one has $\rho_{g}(G_{H})\subset\rho_{g}(G_{H_{0}})\subset\rho_{g}(G_{{\mathbb{Q}}})$, the first inclusion is proper and $[\rho_{g}(G_{{\mathbb{Q}}}):\rho_{g}(G_{H})]=2$, one has $H\neq H_{0}$ and $\rho_{g}(G_{{\mathbb{Q}}})=\rho_{g}(G_{H_{0}})$.

### v3

@@ 4 Examples and counter-examples
Thus either $\alpha u,\alpha u^{-1}$ both lie in ${\mathbb{Z}}_{p}^{\times}$, or they are Galois conjugates of each other lying in some quadratic extension of ${\mathbb{Q}}_{p}$. Since $p$ is odd and prime to the order of $u$, and since $\alpha^{2}\in{\mathbb{Z}}_{p}^{\times}$, $\alpha u,\alpha u^{-1}$ are contained in the maximal unramified extension of ${\mathbb{Q}}_{p}$, so the absolute arithmetic Frobenius $\varphi$ of ${\mathbb{Q}}_{p}^{unr}$ either fixes them both or exchanges them. When $\alpha\notin{\mathbb{Q}}_{p}$, since $\alpha^{2}\in{\mathbb{Q}}$, one has $\varphi(\alpha)=-\alpha$, hence $\varphi(\alpha u)=-\alpha u^{p}$. We thus reach a contradiction if $1\notin\{u^{p-1},u^{p+1}\}$. When $\alpha\in{\mathbb{Q}}_{p}$, one has $\varphi(\alpha u)=\alpha u^{p}$, so we reach a contradiction if $1\notin\{u^{p-1},u^{p+1}\}$.
Assuming that $\varepsilon_{f}$ is the trivial character, that $F={\mathbb{Q}}$, and that $\mathfrak{p}$ is coprime to $2|\rho_{g}(G_{{\mathbb{Q}}})|$, we have thus proved the following. Pick, for every coset $C$ of $G_{H(\mu_{p^{\infty}})}$ in $G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$, some $\alpha_{C}\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ given by Corollary 2, and let $V_{C}\subset L_{0}$ be a set containing the eigenvalues of non-scalar matrices $N\in\rho_{g}(C)$ such that $\det{N}\neq 1$. If, for each $C$ such that $\alpha_{C}\notin{\mathbb{Z}}_{p}^{\times}$, $V_{C}$ does not contain an element $u$ such that $1\in\{u^{p-1},u^{p+1}\}$, and, for each $C$ such that $\alpha_{C}\in{\mathbb{Z}}_{p}^{\times}$, $V_{C}$ does not contain an element such that $1\in\{u^{p-1},u^{p+1}\}$, then $T_{f,g,\mathfrak{p}}$ does not satisfy (wE).
@@
Let $D$ be the derived subgroup of $\rho_{g}(G_{{\mathbb{Q}}})$ and $Z\leq\rho_{g}(G_{{\mathbb{Q}}})$ be the subgroup of scalar matrices. Since $g$ is of $A_{5}$-type, $D\subset\operatorname{SL}_{2}(L_{0})$ is a finite group whose projective image is $A_{5}$. In particular, it has even order, hence contains $I_{2}$. Thus we have an exact sequence $1\rightarrow\{\pm I_{2}\}\rightarrow D\rightarrow A_{5}\rightarrow 1$. Moreover, $D$ and $\rho_{g}(G_{{\mathbb{Q}}})$ have the same projective image, so $\rho_{g}(G_{{\mathbb{Q}}})=ZD$. Since $\det{\rho_{g}(G_{{\mathbb{Q}}})}=\{\pm 1\}$ and $D\subset\operatorname{SL}_{2}(L_{0})$, one has $Z=\{\pm iI_{2},\pm I_{2}\}$, so that $\rho_{g}(G_{{\mathbb{Q}}})=D\cup iD$. Since $D\subset\rho_{g}(G_{H})\subset\rho_{g}(G_{{\mathbb{Q}}})$ and $\varepsilon_{g}(G_{H})=1$, one has $\rho_{g}(G_{H})=D$, and $\rho_{g}(G_{{\mathbb{Q}}}\backslash G_{H})=iD$.
[…]
    if $\tilde{N}$ has order $2$, then $N^{2}$ is scalar, so Cayley–Hamilton implies that $N^{2}=I_{2}$, so the eigenvalues of $N$ are $\pm 1$, so $u^{p\pm 1}=1\neq-1$.
    if $\tilde{N}$ has order $3$, then $N^{3}$ is scalar. Since $\det{N}=-1$, the eigenvalues of $N$ are $\{ij,ij^{2}\}$ or $\{-ij,-ij^{2}\}$. Either way, any eigenvalue $u$ of $N$ has order $12$, and $p\equiv\pm 1\pmod{12}$, so $1\notin\{u^{p-1},u^{p+1}\}$.
    if $\tilde{N}$ has order $5$, then $N^{5}$ is scalar. Since $\det{N}=-1$, any eigenvalue $u$ of $N$ is a root of unity of order $20$. Since $p\not\equiv\pm 1\pmod{5}$, the order of $u^{p\pm 1}$ is divisible by $5$, so $1\notin\{u^{p-1},u^{p+1}\}$ and we are done.
@@
Let $D=\rho_{g}(G_{{\mathbb{Q}}})^{\prime}$ and $Z$ denote the group of scalar matrices contained in $\rho_{g}(G_{{\mathbb{Q}}})$. One has $\det{Z}\subset\varepsilon_{g}(G_{{\mathbb{Q}}})=\{\pm 1\}$ so $Z\subset\{\pm I_{2},\pm iI_{2}\}$. Moreover, the coefficient field of $g$ is ${\mathbb{Q}}(\sqrt{-2})$, so that $g$ has no coefficient equal to $\pm 2i$, hence $Z\subset\{\pm I_{2}\}$. As in the previous case, $D\subset\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})$ and its projective image is $S_{4}^{\prime}=A_{4}$, so $D$ has even order, hence $I_{2}\in D$ and $Z=Z\cap D=\{\pm I_{2}\}$. Thus $D$ is a subgroup of $\rho_{g}(G_{{\mathbb{Q}}})$ with index $2$; since $\varepsilon_{g}(G_{H})=\{1\}$, one furthermore has $D\subset\rho_{g}(G_{H})\subset\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})$: thus $D=\rho_{g}(G_{H})=\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})$ and $\rho_{g}(G_{{\mathbb{Q}}}\backslash G_{H})=\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{GL}_{2}(L_{0})_{\det=-1}$.
[…]
Because $p\equiv\pm 1\pmod{8}$ and $p\not\equiv\pm 1\pmod{12}$, $2$ is a square in ${\mathbb{Z}}_{p}^{\times}$ but $3$ is not, so $\alpha\notin{\mathbb{Z}}_{p}^{\times}$. Thus, we need to prove that 1-1 is not contained in $\{1^{p\pm 1},(-1)^{p\pm 1},\omega^{p\pm 1}\}=\{1,\omega^{p\pm 1}\}$ where $\omega$ is a primitive $8$-th root of unity. Since $p\equiv\pm 1\pmod{8}$, $p\pm 1$ is congruent modulo $8$ to 0,2,20,2,-2, so $\{\omega^{p\pm 1}\}\subset\{1,\omega^{2},\omega^{-2}\}=\{1,\pm i\}$ does not contain 1-1, and we are done.
(c) According to the LMFDB, $L,F,H$ and $\varepsilon_{f},\varepsilon_{g}$ are as above (so $\varepsilon_{g}$ is the quadratic character of $H/{\mathbb{Q}}$), and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\simeq A_{4}$, while $\rho_{g}(G_{{\mathbb{Q}}})$ has cardinality $48$: hence the scalar matrices contained in $\rho_{g}(G_{{\mathbb{Q}}})$ are exactly $\{\pm iI_{2},\pm I_{2}\}$. Now, $\rho_{g}(G_{H})$ is a subgroup of index at most $2$ of $\rho_{g}(G_{{\mathbb{Q}}})$; since it is contained in $\operatorname{SL}_{2}(L_{0})$, it does not contain $iI_{2}$, so $\rho_{g}(G_{{\mathbb{Q}}}\backslash G_{H})=i\rho_{g}(G_{H})$. In particular, $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$.
[…]
Hence we need to show that for any non-scalar matrix $M\in\rho_{g}(G_{{\mathbb{Q}}}\backslash G_{H})=i\rho_{g}(G_{H})$, for any eigenvalue $u$ of $M$, one has $1\notin\{u^{p-1},u^{p+1}\}$.
@@
    If $\tilde{N}$ has order $3$, then as in case (a), any eigenvalue $u$ of $N$ has order $12$. Since $p\equiv\pm 1\pmod{12}$, one among $u^{p-1}$ and $u^{p+1}$ is equal to $1$, and the other has order $6$, so neither equals 1-1 and we are done.
@@
[…]
@@
    If $\det{N}=-1$ and the projective image of $N$ has order $2$, its eigenvalues are $\pm 1$; one has $1^{\pm 1}=(-1)^{p\pm 1}=1\neq-1$, so we are done.
    If $\det{N}=-1$ and the projective image of $N$ has order $6$, and $u$ is an eigenvalue of $N$, then $N\sim\Delta_{u,-u^{-1}}$, $u^{6}=u^{-6}$, $u^{3}\neq-u^{-3}$, $u^{2}\neq u^{-2}$. Hence $u^{2}\neq 1$, $u^{6}=\pm 1$ and $u^{6}\neq-1$, so $u^{6}=1$. Therefore, $u^{3(p\pm 1)}=1$, so $1\notin\{u^{p\pm 1}\}$.
(e) In this case, $L={\mathbb{Q}}(\sqrt{7}),F={\mathbb{Q}}$ and $H={\mathbb{Q}}(\sqrt{-3})$; $g$ has complex multiplication by ${\mathbb{Q}}(\sqrt{-51})$, $\varepsilon_{g}$ is a character of conductor $51$ and order $4$, and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\simeq S_{3}$; $\varepsilon_{f}$ is the trivial character. Since $\varepsilon_{g}(2893)=\pm i$, $\varepsilon_{g}(G_{H})=\{\pm 1,\pm i\}=\varepsilon_{g}(G_{{\mathbb{Q}}})$. The LMFDB shows that $\rho_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $S_{3}\times{\mathbb{Z}}/8{\mathbb{Z}}$. Its center (which is scalar, by Schur’s lemma) is thus the set of $8$-th roots of unity, and $\rho_{g}(G_{{\mathbb{Q}}})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is onto with kernel $\{\pm I_{2}\}$.
The LMFDB data stipulates that $\ker{\tilde{\rho}_{g}}\not\subset G_{H}$, so $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$. Using MAGMA [1], we can also see that $\ker{\rho_{g}}\subset G_{H}$, so that $\rho_{g}(G_{{\mathbb{Q}}}):\rho_{g}(G_{H})]=2$.
@@
Since $\rho_{g}(G_{H})$ and $\rho_{g}(G_{{\mathbb{Q}}})$ have the same determinant and projective image, we can split elements of $\rho_{g}(G_{{\mathbb{Q}}})$ in six categories: $\{\pm iI_{2},\pm I_{2}\}$, scalar matrices of order $8$, matrices in $C$ with determinant $\pm 1$, matrices in $C$ with determinant $\pm i$, matrices outside $C$ with determinant $\pm 1$, matrices outside $C$ with determinant $\pm i$. Then $\rho_{g}(G_{H})$ is exactly made with the matrices of the first, third, and last categories.
[…]
@@
    If $u$ is any eigenvalue of any matrix $N$ in the third or last category such that $\det{N}\neq 1$, then $p$ is not congruent to $\pm 1$ modulo the order of $u$.
    If $u$ is any eigenvalue of any matrix $N$ in the fourth or fifth category such that $\det{N}\neq 1$, then $1\notin\{u^{p-1},u^{p+1}\}$.
Let $N\in C$ be in the third category with $\det{N}\neq 1$. Then $\det{N}=-1$ and $N^{3}$ is scalar with determinant 1-1, so as in (a) any eigenvalue $u$ of $N$ has order $12$. We are done since $p\not\equiv\pm 1\pmod{12}$.
@@
Let $N\in C$ be a matrix in the fourth category. Then $\det{N}=\pm i$ and $N^{3}$ is scalar. Then $N\sim\omega\Delta_{j,j^{2}}$ for some primitive $8$-th root of unity $\omega$, so any eigenvalue $u$ of $N$ has order $24$. Since $p\not\equiv\pm 1\pmod{12}$, none of $u^{2(p\pm 1)}\neq 1$ so $1\notin\{u^{p\pm 1}\}$, and we are done.
Let $N$ be a matrix in the fifth category such that $\det{N}\neq 1$. Then $\det{N}=-1$ and the projective image of $N$ has order $2$, so the eigenvalues of $N$ are $\pm 1$, $1^{p\pm 1}=(-1)^{p\pm 1}=1$, so we are done.
@@
The proposition below, on the other hand, shows that, even in the case where the assumption of Proposition 2 is satisfied (and therefore the reasoning of Corollary 2 does not apply), the representation can still have Euler-adapted Galois image.
@@
    $\rho(G_{{\mathbb{Q}}}):\rho(G_{H})]=[\tilde{\rho}_{g}(G_{{\mathbb{Q}}}):\tilde{\rho}_{g}(G_{H})]=2$,
@@
Proof. –  If $\rho_{g}(G_{H})$ contains a matrix with eigenvalues $1$ and 1-1, we are done by Corollary 2. We may thus assume that $\rho_{g}(G_{H})$ does not contain any matrix whose eigenvalues are $1$ and 1-1.
[…]
@@
Since $G_{H}/G_{HK}$ is a group of order $2$, it follows that $\varepsilon_{g}(G_{HK})=\{\pm 1\}$, so $\varepsilon_{g}(G_{H}\backslash G_{K})=\{\pm i\}$. Let $H_{0}\subset H$ be the subfield such that $\ker{\varepsilon_{f}}=G_{H_{0}}$. One has $\rho_{g}(G_{H})\subset\rho_{g}(G_{H_{0}})\subset\rho_{g}(G_{{\mathbb{Q}}})$, the first inclusion is proper and $\rho_{g}(G_{{\mathbb{Q}}}):\rho_{g}(G_{H})]=2$ by assumption, so $H\neq H_{0}$ and $\rho_{g}(G_{{\mathbb{Q}}})=\rho_{g}(G_{H_{0}})$. In particular, $\rho_{g}(G_{H_{0}})$ contains the image $N_{-}$ of the complex conjugation, which is not contained in $\rho_{g}(G_{H})$ by assumption, since its eigenvalues are 1,11,-1.